11 February 2008

RHIC pp Set-Up:  BBQ blew up emittances.  15 minutes with BBQ on caused emittance to blow up from 13pi to 28pi, due to increased voltages for BTF.

Most of the day yesterday was spent working on rotator ramp for pprot83. 

Power dip brought down the AGS which gave trouble most of the night. 

Provided collisions overnight without rotators, although intensities were low 0.7-0.8e11 in both.

Two Quench Link Interlocks overnight.    One was due to a fibre optic interface card failure, and the other was due to a Yellow Main PS. 

JET is still down due to a failure of the Breit-Rabi polarimeter gauge.  Need 4 hour access.

So today we need to continue rotator ramp development.  Nonlinear chromaticity correction. 


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020